What is the lagrange's equation of simple pendulum?


1
Expert's answer
2021-02-07T19:09:40-0500

Let the length of the pendulum be l, then we can represent the co-ordinates of the motion of the pendulum as per the given question


"x(\theta)= l\sin\theta"

"y(\theta) = -l\cos\theta"

now, differentiating with respect to t,

"\frac{d\theta}{dt}=\dot{\theta}"


"\frac{d(x(\theta))}{dt}=l\cos\theta \frac{d\theta}{dt}"


"\Rightarrow \frac{d(x(\theta))}{dt}=l\dot{\theta} \cos\theta"


"\frac{d(y(\theta))}{dt}=l\sin\theta \frac{d\theta}{dt}"

"\Rightarrow \frac{d(y(\theta))}{dt}=l\dot{\theta}\sin\theta"

Now, applying the conservation of energy

"L=T-V=KE-PE"


"=\frac{ml\dot{\theta^2}}{2}-mgl(1-\cos\theta)"


We know that the Lagrange equation for the generalized co-ordinate

"\frac{d}{dt}(\frac{(\partial L)}{d\dot{\theta}})-\frac{\partial L}{\partial{\theta}}=0"


"\Rightarrow \ddot{\theta}+\frac{g}{l}\sin(\theta)=0"

Hence, natural frequency of the simple pendulum

"\omega = \sqrt{\frac{g}{l}}"

"T=2\pi \sqrt{\frac{l}{g}}"